# On the pro-modularity in the residually reducible case for some totally real fields

### v1

@@ 3. Galois deformation theory
3.1. Properties of universal deformation rings
@@
Let $R^{\textnormal{ps}}$ be the universal pseudo-deformation ring of $1+\bar{\chi}$, and $T^{\textnormal{univ}}$ the universal deformation. The existence of $(R^{\textnormal{ps}},T^{\textnormal{univ}})$ is proved in [SW99, Section 2.4].
@@
Let $\delta_{F}$ be the Leopoldt defect (see [NSW13, 10.3.7].
1) $\operatorname{dim}(R_{x}^{\textnormal{aux,red}})\leq 1+\delta_{F}+\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1})))$.
2) $\operatorname{dim}(R_{x}^{\textnormal{aux,red}})\geq 2[F:\mathbb{Q}]+1-\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi})))$.
@@
The proof of the first part is an analogue to [SW99, Section 2.2]. We sketch a proof here.
@@
The second part is a consequence of the results of GMA. We choose a basis for $\rho_{x}^{\textnormal{aux}}$ such that $\rho_{x}^{\textnormal{aux}}(c)=\begin{pmatrix}1&~{}\\ ~{}&-1\end{pmatrix}$ and write $\rho_{x}^{\textnormal{aux}}(\sigma)=\begin{pmatrix}a_{\sigma}&b_{\sigma}\\ c_{\sigma}&d_{\sigma}\end{pmatrix}$. Then by Proposition 2.2.3, $A=R_{x}^{\textnormal{aux}}[\rho_{x}^{\textnormal{aux}}(\operatorname{Gal}(F_{\Sigma}/F))]$ is a GMA over $R_{x}^{\textnormal{aux}}$, a sub-$R_{x}^{\textnormal{aux}}$-GMA of $M_{2}(R_{x}^{\textnormal{aux}})$.
Now write $A=\begin{pmatrix}R_{x}^{\textnormal{aux}}&B\\ C&R_{x}^{\textnormal{aux}}\end{pmatrix}$. Then $C$ is the ideal of $R_{x}^{\textnormal{aux}}$ generated by all $c_{\sigma}$. By Proposition 2.2.4 and Nakayama’s lemma, $C$ can be generated by at most $\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi})))$ elements as $\operatorname{Ext}^{1}(\bar{\chi},1)\cong H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}))$. By the universal property, we also have $R_{x}^{\textnormal{aux,red}}\cong R_{x}^{\textnormal{aux}}/C$. Combining Proposition 3.1.2, we know that $\operatorname{dim}(R_{x}^{\textnormal{aux,red}})\geq 2[F:\mathbb{Q}]+1-\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi})))$.
Proposition 3.1.5.
Assume that $F$ is an abelian totally real field. Further assume that $[F:\mathbb{Q}]=\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1})))=\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi})))$. Then $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$.
Proof.
From our assumptions and Proposition 3.1.4, the Krull dimension of $R_{x}^{\textnormal{aux,red}}$ is $[F:\mathbb{Q}]+1$. By 1) of Proposition 3.1.2, we know that $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$. ∎
Remark 3.1.6.
1) This result is a generalization of [Deo23, Corollary 3.5] to abelian totally real fields. In Deo’s result, he also assumes the equation of the dimension $1(=[F:\mathbb{Q}])=\operatorname{dim}_{\mathbb{F}}(H^{1}(\mathbb{Q}_{\Sigma}/\mathbb{Q},\mathbb{F}(\bar{\chi}^{-1})))$, and in this case, we can get a surjection from the universal pseudo-deformation ring $R^{\textnormal{ps}}$ to the universal deformation ring $R_{x}$ by Kisin’s argument [Kis09, Corollary 1.4.4] which is a crucial step in Deo’s proof. However, we cannot assume that the cohomology group $H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$ is of dimension $1$ in the totally real field case because the global characteristic formula tells us that it is of dimension at least $[F:\mathbb{Q}]$. Thus, we have to give a new method to prove it.
2) The condition in our result can be studied by Galois cohomology. See the appendix.
3.2. Some miscellaneous results
1) $\chi$ is unramified outside $\Sigma_{p}$,
@@ 3.2. Some miscellaneous results
Let $(R_{\textnormal{aux}}^{\textnormal{ps}},T_{\textnormal{aux}})$ be the universal pseudo-deformation of $\operatorname{tr}(\bar{\rho}_{0})$ of auxiliary type (i.e., with $\operatorname{det}=\chi$) and $(R^{\textnormal{ps},\{\xi_{v}\}},T^{\textnormal{ps},\{\xi_{v}\}})$ the universal pseudo-deformation of $\operatorname{tr}(\bar{\rho}_{0})$ of auxiliary type satifying $T_{\textnormal{aux}}\mid_{I_{v}}=\xi_{v}+\xi_{v}^{-1}$ for all $v\in\Sigma\setminus\Sigma_{p}$. Clearly, there are natural surjections $R^{\textnormal{ps}}\twoheadrightarrow R_{\textnormal{aux}}^{\textnormal{ps}}\twoheadrightarrow R^{\textnormal{ps},\{\xi_{v}\}}$. By [Pan22, Corollary 5.7.2], for any minimal prime $\mathfrak{P}$ of $R_{\textnormal{aux}}^{\textnormal{ps}}$, the natural surjection $R_{\textnormal{aux}}^{\textnormal{ps}}\to R_{\textnormal{aux}}^{\textnormal{ps}}/\mathfrak{P}$ factors through some $R^{\textnormal{ps},\{\xi_{v}\}}$. Similarly, we also say that $\mathfrak{P}$ belongs to $\{\xi_{v}\}$ in this situation.
@@
Let $(R^{\textnormal{ps,red}},T^{\textnormal{red}})$ be the universal pseudo-deformation ring of auxiliary type such that $T^{\textnormal{red}}$ is reducible (i.e. $y(\sigma,\tau)=0$ for all $\sigma,\tau\in\operatorname{Gal}(F_{\Sigma}/F)$). The following proposition is just an analogue to [SW99, Lemma 2.9].
@@
We have $\operatorname{dim}(R^{\textnormal{ps,red}})\leq 2+\delta_{F}$.
@@
Consider the character $d:\operatorname{Gal}(F_{\Sigma}/F)\to(R^{\textnormal{ps,red}})^{\times}$ in the sense of Definition 2.1.1, and it induces a natural surjection $\mathcal{O}[[\operatorname{Gal}(L(\Sigma)/F)]]\twoheadrightarrow R^{\textnormal{ps,red}}$. Hence, $\operatorname{dim}(R^{\textnormal{ps,red}})\leq 2+\delta_{F}$. ∎
@@
Suppose that $\mathfrak{s}$ is a prime of $R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $\pi$ of dimension at least $2$. If $F$ is abelian over $\mathbb{Q}$, then we can find a one-dimensional prime $\mathfrak{r}$ of $R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $\mathfrak{s}$ such that $T_{\textnormal{aux}}~{}\operatorname{mod}~{}\mathfrak{r}$ is irreducible.
@@
For simplicity, we suppose that $R=R_{\textnormal{aux}}^{\textnormal{ps}}/\mathfrak{s}$ is of dimension $2$. For any one-dimensional prime $\mathfrak{t}$ of $R$ such that $T_{\textnormal{aux}}~{}\operatorname{mod}~{}\mathfrak{t}$ is reducible, we get a surjection $\mathcal{O}[[\operatorname{Gal}(L(\Sigma)/F)]]/(\pi)\twoheadrightarrow R^{\textnormal{ps,red}}/(\pi)\twoheadrightarrow R/\mathfrak{t}$, all of which have dimension $1$ by Proposition 3.2.3. Consider an element $t\in\mathfrak{m}_{R}$ but $t\notin\mathfrak{t}$, and let $\sqrt{(t)}=\cap\mathfrak{t}_{i}$ be the primary decomposition. If $\mathfrak{t}_{i}$ is one-dimensional for some $i$, then we can take $\mathfrak{r}=\mathfrak{t}_{i}$. Otherwise, we get $\sqrt{(t)}=\mathfrak{m}_{R}$ as $\operatorname{dim}R=2$. Then by Krull’s principal ideal theorem, we have $10=\operatorname{dim}(R/tR)\geq\operatorname{dim}R-1=$ 1, which is impossible. ∎
@@ 4. Pro-modularity
In this section, we will prove the pro-modularity of the universal deformation ring. We fix a totally real field $F$ in which $p>3$ splits completely.
@@ 4.2. Deformation rings and Hecke algebras
Let $F$ be an abelian totally real number field of even degree over $\mathbb{Q}$ in which $p$ splits completely. As in Section 3.2, we assume that for all $v\in\Sigma\setminus\Sigma_{p}$, we have $p\mid\operatorname{Nm}(v)-1$. Let $\xi_{v}:k(v)^{\times}\to\mathcal{O}^{\times}$ be characters of $p$-power order for $v\in\Sigma\setminus\Sigma_{p}$, and we can view them as characters of $I_{v}$ by local class field theory. Let $\chi:\operatorname{Gal}(F_{\Sigma}/F)\to\mathcal{O}^{\times}$ be a continuous totally odd character such that
@@
3) If $\rho(\mathfrak{q})$ is dihedral, namely isomorphic to $\operatorname{Ind}_{G_{L}}^{G_{F}}\theta$ for some quadratic extension $L$ of $F$ and continuous character $\theta:G_{L}\to k(\mathfrak{q})^{\times}$, then $L\cap F(\zeta_{p})=F$, where $\zeta_{p}$ is a primitive $p$-th root of unity.
@@
Now we consider the deformation rings and Hecke algebras in the ordinary case.
@@
Now we define the Iwasawa algebra
@@
and we get a natural map $\Lambda_{F}\to R^{\textnormal{ps,ord}}$.
@@
Corollary 4.2.7.
The Krull dimension of $R^{\textnormal{ps,ord}}$ is $1+[F:\mathbb{Q}]$.
@@
This result is based on [Pan22, Section 5]. We just sketch a proof here.
By Proposition 4.2.6, we know that $\operatorname{dim}R^{\textnormal{ps,ord}}\leq 1+[F:\mathbb{Q}]$. If we consider the ordinary Hecke algebra $\mathbb{T}^{\textnormal{ord}}$(see [Pan22, Section 5.2] for the explicit definition), by [Pan22, Corollary 5.2.1], it is equidimensional of dimension $1+[F:\mathbb{Q}]$. As there is a two-dimensional pseudo-representation $T^{\textnormal{ord}}:\operatorname{Gal}(F_{\Sigma}/F)\to\mathbb{T}^{\textnormal{ord}}$, we get a natural surjection $R^{\textnormal{ps,ord}}\twoheadrightarrow\mathbb{T}^{\textnormal{ord}}$, and we have $\operatorname{dim}R^{\textnormal{ps,ord}}\geq 1+[F:\mathbb{Q}]$. Hence, $\operatorname{dim}R^{\textnormal{ps,ord}}=1+[F:\mathbb{Q}]$. ∎
Definition 4.2.8.
Proposition 4.2.9.
Proof.
See [Pan22, Lemma 5.3.2]. ∎

### v3

@@ 3. Galois deformation theory
3.1. Universal deformation rings
@@
Let $\delta_{F}$ be the Leopoldt defect (see [NSW13, 10.3.7]. Then we have $\operatorname{dim}(R_{x}^{\textnormal{aux,red}})\leq 1+\delta_{F}+\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1})))$.
@@
The proof is an analogue to [SW99, Section 2.2]. We sketch a proof here.
@@
3.2. Universal pseudo-deformation ring
1) $\chi$ is unramified outside $\Sigma_{p}$,
Let $R^{\textnormal{ps}}$ be the universal pseudo-deformation ring of $\operatorname{tr}(\bar{\rho}_{0})=1+\bar{\chi}$, and $T^{\textnormal{univ}}$ the universal deformation. The existence of $(R^{\textnormal{ps}},T^{\textnormal{univ}})$ is proved in [SW99, Section 2.4]. Let $(R_{\textnormal{aux}}^{\textnormal{ps}},T_{\textnormal{aux}})$ be the universal pseudo-deformation of $\operatorname{tr}(\bar{\rho}_{0})$ of auxiliary type (i.e., with $\operatorname{det}=\chi$) and $(R^{\textnormal{ps},\{\xi_{v}\}},T^{\textnormal{ps},\{\xi_{v}\}})$ the universal pseudo-deformation of $\operatorname{tr}(\bar{\rho}_{0})$ of auxiliary type satifying $T_{\textnormal{aux}}\mid_{I_{v}}=\xi_{v}+\xi_{v}^{-1}$ for all $v\in\Sigma\setminus\Sigma_{p}$. Clearly, there are natural surjections $R^{\textnormal{ps}}\twoheadrightarrow R_{\textnormal{aux}}^{\textnormal{ps}}\twoheadrightarrow R^{\textnormal{ps},\{\xi_{v}\}}$. By [Pan22, Corollary 5.7.2], for any minimal prime $\mathfrak{P}$ of $R_{\textnormal{aux}}^{\textnormal{ps}}$, the natural surjection $R_{\textnormal{aux}}^{\textnormal{ps}}\to R_{\textnormal{aux}}^{\textnormal{ps}}/\mathfrak{P}$ factors through some $R^{\textnormal{ps},\{\xi_{v}\}}$. Similarly, we also say that $\mathfrak{P}$ belongs to $\{\xi_{v}\}$ in this situation.
@@ 3.2. Universal pseudo-deformation ring
Let $(R^{\textnormal{ps,red}},T^{\textnormal{red}})$ be the universal pseudo-deformation ring of auxiliary type such that $T^{\textnormal{red}}$ is reducible, i.e., $y(\sigma,\tau)=0$ for all $\sigma,\tau\in\operatorname{Gal}(F_{\Sigma}/F)$.
@@
Assume $F$ is abelian over $\mathbb{Q}$.
1) We have $\operatorname{dim}(R^{\textnormal{ps,red}})\leq 2$.
2) Assume $\bar{\chi}$ is quadratic and can be extended to $G_{\mathbb{Q}}$. Then the dihedral locus $C^{\operatorname{dih}}$ of $\operatorname{Spec}R_{\textnormal{aux}}^{\textnormal{ps}}$, i.e., for any point $\mathfrak{p}$ in $C^{\operatorname{dih}}$, $\rho(\mathfrak{p})\cong\operatorname{Ind}_{G_{L}}^{G_{F}}\theta$ for some character $\theta$ and the splitting field $L=F(\bar{\chi})$, is of dimension at most $2+[F:\mathbb{Q}]$.
@@
1) Consider the character $d:\operatorname{Gal}(F_{\Sigma}/F)\to(R^{\textnormal{ps,red}})^{\times}$ in the sense of Definition 2.1.1, and it induces a natural surjection $\mathcal{O}[[G_{F}^{ab}]]\twoheadrightarrow R^{\textnormal{ps,red}}$. Hence, $\operatorname{dim}(R^{\textnormal{ps,red}})\leq 2+\delta_{F}=2$. The last equation is from Leopoldt’s conjecture for abelian totally real fields.
2) From our definition, we can find that if $\mathfrak{p}\in C^{\operatorname{dih}}$ and $\rho(\mathfrak{p})\cong\operatorname{Ind}_{G_{L}}^{G_{F}}\theta$, then the global character $\theta$ is a lifting of the character $\bar{\chi}|_{G_{L}}=\mathbf{1}$. Thus, the dihedral locus is at most of $\operatorname{dim}\mathcal{O}[[G_{L}^{ab}]]$, which is $2+[F:\mathbb{Q}]$ by Leopoldt’s conjecture for abelian CM fields. ∎
@@
Keep assumptions as in Proposition 3.2.2. Let $\mathfrak{s}$ be a prime of $R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $p$ and write $R=R_{\textnormal{aux}}^{\textnormal{ps}}/\mathfrak{s}$.
1) If $\dim R\geq 2$, then we can find a prime $\mathfrak{s}^{\prime}$ of $R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $\mathfrak{s}$ such that the corresponding representation is irreducible.
2) If $\dim R\geq 2+[F:\mathbb{Q}]$, then we can find a prime $\mathfrak{s}^{\prime}$ of $R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $\mathfrak{s}$ such that the corresponding representation is irreducible and not dihedral.
@@
We first give a useful lemma in commutative algebra.
Lemma 3.2.4.
Let $R$ be a noetherian local ring with maximal ideal $\mathfrak{m}_{R}$. Suppose that $V\subsetneqq\operatorname{Spec}R$ is a closed subset. Then there exists a one-dimensional prime $\mathfrak{p}\notin V$.
Proof of Lemma 3.2.4.
Using [Sta23, Lemma 28.6.4], the scheme $\operatorname{Spec}R\setminus\{\mathfrak{m}_{R}\}$ is Jacobson, i.e., the underlying topological space is Jacobson. As $V\neq\operatorname{Spec}R$, there exists a closed point of $\operatorname{Spec}R\setminus\{\mathfrak{m}_{R}\}$ not contained in $V$. This shows the result. ∎
Our assertions are direct consequences from Lemma 3.2.4 and Proposition 3.2.2. ∎
@@
Definition 3.2.6.
[Pan22, Definition 5.4.3] Let $R$ be a CNL ring. The connectedness dimension of $R$ is defined to be
$c(R):=\min_{Z_{1},Z_{2}}\{\dim(Z_{1}\cap Z_{2})\},$
where $Z_{1},Z_{2}$ are non-empty unions of irreducible components of $\operatorname{Spec}R$ such that $Z_{1}\cup Z_{2}=\operatorname{Spec}R$.
Proposition 3.2.7.
Suppose $\mathfrak{p}$ is a one-dimensional prime of $R_{\textnormal{aux}}^{\textnormal{ps}}$ whose corresponding represenation is absolutely irreducible.
1) $\operatorname{dim}(R_{\textnormal{aux}}^{\textnormal{ps}})_{\mathfrak{p}}\geq 2[F:\mathbb{Q}]$.
2) The connectedness dimension $c((R_{\textnormal{aux}}^{\textnormal{ps}})_{\mathfrak{p}})$ is at least $2[F:\mathbb{Q}]-1$.
Proof.
For the first part, see [Pan22, Lemma 7.1.2]. For the second part, see [Pan22, Lemma 7.4.6]. ∎
Remark 3.2.8.
Unlike the universal deformation ring $R_{x}^{\textnormal{aux}}$, the structure of the universal pseudo-deformation ring $R_{\textnormal{aux}}^{\textnormal{ps}}$ is mysterious because we cannot use the global characteristic formula as in Proposition 3.1.2. However, we can show that if $\mathfrak{p}$ is a one-dimensional irreducible prime of $R_{\textnormal{aux}}^{\textnormal{ps}}$, then every irreducible component of $\operatorname{Spec}R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $\mathfrak{p}$ is of dimension at least $2[F:\mathbb{Q}]+1$, as an analogue to Proposition 3.1.3. The proof is as follows.
If our claim is not true, then we have a partition $Z_{1}\amalg Z_{2}=\operatorname{Spec}(R_{\textnormal{aux}}^{\textnormal{ps}})_{\mathfrak{p}}$, where $Z_{i}$ are finite sets of some irreducible components of $\operatorname{Spec}(R_{\textnormal{aux}}^{\textnormal{ps}})_{\mathfrak{p}}$ such that $Z_{1}$ consists of all ones of dimension at least $2[F:\mathbb{Q}]$. By the first part of Proposition 3.2.7 and our assumption, both $Z_{1}$ and $Z_{2}$ are non-empty. By the second part of Proposition 3.2.7, we can find $C_{1}\in Z_{1},C_{2}\in Z_{2}$ and a prime $\mathfrak{q}\in C_{1}\cap C_{2}$ of dimension at least $2[F:\mathbb{Q}]-1$. As $\operatorname{dim}C_{2}\leq 2[F:\mathbb{Q}]-1$, we know that $\mathfrak{q}$ is the generic point of $C_{2}$, which is against to the fact $\mathfrak{q}\in C_{1}$. Thus, $Z_{2}$ has to be empty.
We do not know whether every irreducible component of $R_{\textnormal{aux}}^{\textnormal{ps}}$ is of dimension at least $2[F:\mathbb{Q}]+1$.
@@ 4. Pro-modularity
In this section, we will prove the pro-modularity of the universal deformation ring. We fix a totally real field $F$ and a prime $p>3$.
@@ 4.1. Hecke algebras
Assume that the prime $p$ splits completely in the totally real field $F$.
@@ 4.2. Deformation rings and Hecke algebras
Assume $F$ is an abelian totally real number field of even degree over $\mathbb{Q}$ in which $p$ splits completely. As in Section 3.2, we assume that for all $v\in\Sigma\setminus\Sigma_{p}$, we have $p\mid\operatorname{Nm}(v)-1$. Let $\xi_{v}:k(v)^{\times}\to\mathcal{O}^{\times}$ be characters of $p$-power order for $v\in\Sigma\setminus\Sigma_{p}$, and we can view them as characters of $I_{v}$ by local class field theory. Let $\chi:\operatorname{Gal}(F_{\Sigma}/F)\to\mathcal{O}^{\times}$ be a continuous totally odd character such that
@@
Remark 4.2.2.
In some cases, the big Hecke algebra is equidimensional of dimension $1+2[F:\mathbb{Q}]$. See the proof of [Pan22, Theorem 3.6.1] and [Pan22, Remark 3.6.10].
@@
3) (dihedral condition) If $\rho(\mathfrak{q})$ is dihedral, namely isomorphic to $\operatorname{Ind}_{G_{L}}^{G_{F}}\theta$ for some quadratic extension $L$ of $F$ and continuous character $\theta:G_{L}\to k(\mathfrak{q})^{\times}$, then $L\cap F(\zeta_{p})=F$, where $\zeta_{p}$ is a primitive $p$-th root of unity.
@@
We briefly summarize some main steps of the proof of Theorem 4.2.5. Here we follow most notations as in [Pan22, Section 4], and one can find more details there.
Using Taylor’s Ihara avoidance trick as in [Tay08], one can show that $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}$ is equidimensional of dimension $4|P|-1+r+2[F:\mathbb{Q}]$ (see [Pan22, Section 4.2, Lemma 4.8.2]). As $(R^{\{\xi_{v}\}})^{\prime}$ is finite free over $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}$, we have
From Theorem 4.2.1, [Eis13, Corollary 12.5], and discussions above, we can also conclude that for a nice prime $\mathfrak{q}$ of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$, we have $\operatorname{dim}\widehat{(\mathbb{T}_{\xi})_{\mathfrak{q}}}=\operatorname{dim}(\mathbb{T}_{\xi})_{\mathfrak{q}}=2[F:\mathbb{Q}]$. In other words, if an irreducible component of the big Hecke algebra contains a nice prime, then it is of dimension $1+2[F:\mathbb{Q}]$.
Remark 4.2.7.
@@
4.3. The ordinary case
In this subsection, we recall some results on the deformation rings and Hecke algebras in the ordinary case.
Assume that $F$ is an abelian totally real number field over $\mathbb{Q}$ in which $p$ is unramified. Let $\Sigma$ be a finite set of primes of $F$ containing all places above $p$. Let $\chi:\operatorname{Gal}(F_{\Sigma}/F)\to\mathcal{O}^{\times}$ be a continuous totally odd character such that
1) $\bar{\chi}$ can be extended to a character of $G_{\mathbb{Q}}$.
2) $\bar{\chi}\mid_{G_{F_{v}}}$ is not trivial for any $v\mid p$.
3) $\chi\mid_{G_{F_{v}}}$ is de Rham for any $v\mid p$.
@@ 4.3. The ordinary case
Definition 4.3.1.
Proposition 4.3.2.
Proof.
See [Pan22, Lemma 5.3.2]. ∎
Define the Iwasawa algebra
@@
and we get a natural map $\Lambda_{F}\to R^{\textnormal{ps,ord}}$ induced by the characters $\psi_{v,1}^{\textnormal{ord}}$. The following result is the ordinary Fontaine-Mazur conjecture proved in [SW99] and [Pan22, Section 5], which is used to produce pro-modular primes in our setting.
@@
Lemma 4.3.4.
Let $C^{\textnormal{ord}}$ be an irreducible component of $R^{\textnormal{ps,ord}}$ of dimension at least $1+[F:\mathbb{Q}]$. Let $C^{\textnormal{ord,aut}}$ be the set of regular de Rham primes in $C^{\textnormal{ord}}$. Then $C^{\textnormal{ord,aut}}$ is dense in $C^{\textnormal{ord}}$. Here a prime $\mathfrak{q}$ is called regular de Rham prime if
@@
The proof is the same as [Pan22, Corollary 7.2.3]. We sketch it here.
From our assumption and Proposition 4.3.3, we know that the morphism $C^{\textnormal{ord}}\to\operatorname{Spec}\Lambda_{F}$.
For $v\mid p$, let $\psi_{v,1}^{\textnormal{ord}},\psi_{v,2}^{\textnormal{ord}}:G_{F_{v}}\to(R^{\textnormal{ps,ord}})^{\times}$ be the liftings of $1,\bar{\chi}\mid_{G_{F_{v}}}$ respectively. By Proposition 4.3.2, there exists $n_{v}\in\{1,2\}$ such that for any $\mathfrak{q}\in C_{1}^{\textnormal{ord}}$, we get $\rho(\mathfrak{q})\mid_{G_{F_{v}}}\cong\begin{pmatrix}\psi_{v,n_{v}}^{\textnormal{ord}}~{}\operatorname{mod}~{}\mathfrak{q}&*\\ 0&\psi_{v,3-n_{v}}^{\textnormal{ord}}~{}\operatorname{mod}~{}\mathfrak{q}\end{pmatrix}$.
Note that the map $\Lambda_{F}\to R^{\textnormal{ps,ord}}$ is induced by the characters $\psi_{v,1}^{\textnormal{ord}}$. Hence, we know that the image of $C^{\textnormal{ord,aut}}$ is dense in $\operatorname{Spec}\Lambda_{F}$, and hence dense in $C^{\textnormal{ord}}$. ∎